The maximum allowable level of selenium in drinking water is 50ppb. A chemist found 34ug of selenium in 500mL of lake water. Would you drink that lake water? Why or why not?
1
Expert's answer
2020-04-06T11:44:34-0400

The maximum allowable level of selenium in drinking water is 50 ppb, i.e. 50 molecules of selenium per 109 molecules of water.

In 34 ug of Se there is 34*10-6 g / 79 g/mole = 4.3*10-7 mole of Se, or 4.3*10-7*6.02*1023 = 2.59*1017 molecules of selenium.

In 500 mL of water (density = 1 g/mL) there is 500 g / 18 g/mole = 27.78 mole of H2O, or 27.78*6.02*1023 = 1.67*1025 molecules.

The relative amounts of molecules of Se : H2O = 2.59*1017 : 1.67*1025 = 1.55*1017 : 1025 = 15.5 : 109 , i.e. 15.5 ppb of selenium. The water is drinkable, since selenium content is below the allowable limit.